OpenStudy (anonymous):

please help! how can i simplify this fraction 5f+6-f+10 --------- (f-1)(f+8)

OpenStudy (mertsj):

In the numerator add 5f + 1f Also add 6 + 10

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

4f+16

OpenStudy (mertsj):

So now we have: \[\frac{6f+16}{(f-1)(f+8)}\]

OpenStudy (mertsj):

Now pull 2 out of 6f+16

OpenStudy (mertsj):

Please note that 1f+5f=6f NOT 4f

OpenStudy (mertsj):

What do you get when you pull the 2 out?

OpenStudy (anonymous):

wait, but isnt it 5f-1f?

OpenStudy (mertsj):

Oh my. you are right. I thought it said plus.

OpenStudy (mertsj):

So now pull out 4 not 2

OpenStudy (anonymous):

alright, so what does it mean by pull out?

OpenStudy (mertsj):

\[\frac{4(f+4)}{(f-1)(f+8)}\]

OpenStudy (mertsj):

It just means factor out the common factor.

OpenStudy (mertsj):

And now we see that there are no common factors so the fraction cannot be reduced so it is in simple form.

OpenStudy (anonymous):

oh right, okay. so can i simplify it any more?

OpenStudy (anonymous):

alright! thanks so much!!

OpenStudy (mertsj):

Not the way it is stated. But I think you should double check and make sure the 10 is +10 and not -10. It it were -10, it could reduce. Please check.

OpenStudy (anonymous):

okay. yeah its +10

OpenStudy (mertsj):

ok. Good. We're done. I guess all they wanted you to do was collect the like terms in the numerator.

OpenStudy (anonymous):

yeah. thanks!

OpenStudy (mertsj):

yw